OpenStudy (anonymous):

Find the vertex of the parabola and indicate if it is a minimum or maximum for the function f(x) = 7(x - 7)2 - 6 A.vertex at (-7, 6) is a maximum B. vertex at (-7, 6) is a minimum C. vertex at (7, -6) is a minimum D. vertex at (7, -6) is a maximum

terenzreignz (terenzreignz):

okay, what's the vertex here?

terenzreignz (terenzreignz):

remember, if you see a form y = a(x - h)² + k then the vertex is (h,k)

OpenStudy (anonymous):

THIS IS A PRETEST I DONT KNOW HOW TO DO THIS

terenzreignz (terenzreignz):

Calm down, I just told you how to do it :) Look at your equation and identify which values correspond to h and k. y = a( x - h )² + k

OpenStudy (anonymous):

-7 AND 6 CORRESPOND

terenzreignz (terenzreignz):

But that would mean h = -7 and k = 6?

terenzreignz (terenzreignz):

Which would make your equation y = a ( x - (-7) )² + 6 y = a ( x + 7 )² +6 Which is not your equation, is it? :)

OpenStudy (anonymous):

OK SO IS IT MAX. OR MIN.

terenzreignz (terenzreignz):

First, what's your vertex? (and easy on the caps lock there, mate)

OpenStudy (anonymous):

sorry

terenzreignz (terenzreignz):

So your vertex, what is it?

OpenStudy (anonymous):

i dont know what is my vertex

terenzreignz (terenzreignz):

Well then, look at your equation again y = 7(x - 7)² - 6 The vertex form of the parabola goes like y = a(x - h)² + k Where (h,k) is the vertex. Be extra careful with signs.

OpenStudy (anonymous):

-7 and 6 are my vertex

terenzreignz (terenzreignz):

I told you to be extra careful with signs...

OpenStudy (anonymous):

what did i do wrong

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

-7 and -6

terenzreignz (terenzreignz):

Well you see that in the form y = a ( x - h )² + k you have -h and not +h that should figure into your calculations

OpenStudy (anonymous):

but is it max. or min.

terenzreignz (terenzreignz):

We'll get to that once you know the vertex.

OpenStudy (anonymous):

im kinda on a timed homework test can you give me the answer then explain plz

terenzreignz (terenzreignz):

I prefer explaining first so that you can get the answer on your own. You'll feel better that way :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

where did we stop

terenzreignz (terenzreignz):

y = a( x - h )² + k y = 7( x - 7 )² - 6 We were finding the values of h and k.

OpenStudy (anonymous):

so you want me to solve the problem

terenzreignz (terenzreignz):

yes, but hopefully, YOU want you to solve the problem too :)

OpenStudy (anonymous):

how do you make 6 positve

terenzreignz (terenzreignz):

Sorry?

OpenStudy (anonymous):

at the end of the formula its +k i have a -6

terenzreignz (terenzreignz):

Yeah, so you have +k = -6 So what's k?

OpenStudy (anonymous):

i dont know thats why im telling you

terenzreignz (terenzreignz):

That's already it, k = -6 now what's h?

OpenStudy (anonymous):

-7

terenzreignz (terenzreignz):

You see, there'a difference, you have a -h in the vertex form and not just a normal h or +h

OpenStudy (anonymous):

ok

terenzreignz (terenzreignz):

So, solve for h?

OpenStudy (anonymous):

49

OpenStudy (anonymous):

oh i just notice i made a mistake

OpenStudy (anonymous):

the problem is f(x) = 7(x - 7)^2 - 6